# Globally-Constrained Decentralized Optimization with Variable Coupling

### v3

@@ II-A Examples of Applications
Many real-world engineering problems can be cast into the form of (1). Below, we provide two of such examples.
@@ II-A1 A Multi-Vehicle Formation Stabilization Problem
Consider a multi-vehicle formation stabilization problem. Our goal is to stabilize $n$ vehicles on a plane toward an equilibrium point in a decentralized way. For each vehicle $i$, $s_{i}^{k}=[(p_{i}^{k})^{T},(v_{i}^{k})^{T}]^{T}\in\mathbb{R}^{4}$ is the state and $u_{i}^{k}\in\mathbb{R}^{2}$ is the control input at time $k\geq 0$, where $p_{i}^{k}\in\mathbb{R}^{2}$, $v_{i}^{k}\in\mathbb{R}^{2}$ are the position and the velocity of vehicle $i$, respectively. The system dynamic of each vehicle $i$ is given by
@@
Given the initial states $s_{i}^{0}$ $\forall i=1,\ldots,n$, all the vehicles obtain their control actions through model predictive control (MPC), which requires them to cooperatively solve a constrained convex optimization problem as follows.
Let $N\geq 1$ be the prediction horizon in MPC. The optimization variable associated with each vehicle $i$ is $x_{i}=[(s_{i}^{1})^{T},\ldots,(s_{i}^{N})^{T},(u_{i}^{0})^{T},\ldots,(u_{i}^{N-1})^{T}]^{T}\in\mathbb{R}^{6N}$. Then, the multi-vehicle formation stabilization problem is formulated as the following optimization problem
@@
In (3), each $H_{i}$ is a symmetric positive semidefinite matrix and each $Q_{i}\in\mathbb{R}^{6|\mathcal{N}_{i}|N}$, which characterize the cost of vehicle $i$. The objective in (3) is the sum of tracking cost of each vehicle, where each vehicle $i$’s tracking cost is influenced by the positions of its neighboring vehicles [28, 40]. The first constraint in (3) stands for resource limitations and position/velocity requirements across the whole system. The second constraint represents the dynamics of each vehicle $i$ over the prediction horizon, which is directly derived from (2). The last constraint restricts each $x_{i}$ to some compact convex set $X_{i}$ due to certain system limitations. In (3), $H_{i}$, $Q_{i}$, $P_{i}$, $b_{i}$, $C_{i}$, $b_{i}^{\prime}$, and $X_{i}$ are only accessible to vehicle $i$. Clearly, the distributed MPC problem (3) is a particular form of (1).
@@ II-A2 Task Offloading over Cooperative Mobile Edge Computing Networks
Consider a task offloading problem among $n$ edge servers, which form a connected network. Each edge server receives a large quantity of latency-sensitive tasks from exogenous sources, such as mobile devices. Since the computation resources of each edge server are limited and often distinct, each edge server needs to offload part of the received tasks to neighboring edge servers in order to make full use of computation resources[26]. All the edge servers cooperate to find optimal task offloading decisions by minimizing the total completion time of all the tasks subject to the limitations of channel bandwidth and computation capability.
We label the set of edge servers as $\mathcal{V}=\{1,2,\ldots,n\}$. Let $\mathcal{N}_{i}$ be the neighbor set of edge server $i$ including itself. We define $x_{ii}$ and $x_{ij}$ as the quantity of tasks allocated by edge server $i$ to itself and to its neighboring edge server $j$, respectively. Let $x_{ij}$ $\forall j\in\mathcal{N}_{i}$ constitute the decision vector $x_{i}$ of edge server $i$. Also, let the local objective of edge server $i$ be the task completion time for handling its offloaded task, which, apparently, is determined by the tasks offloaded by both edge server $i$ itself and its neighbors, i.e., $x_{ji}$ $\forall j\in\mathcal{N}_{i}$. Thus, we can write the local objective of edge server $i$ as $f_{i}\left(x_{\mathcal{N}_{i}}\right)$ for simplicity. The coupled inequality constraints $\sum_{i\in\mathcal{V}}g_{i}\left(x_{\mathcal{N}_{i}}\right)\leq\mathbf{0}_{p}$ are often associated with the total channel bandwidth or computation resource limit across the network, where $g_{i}$ represents the edge server $i$’s channel bandwidth or computation resource consumption [41]. For each edge server, the sum of the task offloaded to itself and neighboring servers should be equal to the tasks it received from the exogenous sources, which can be described by linear equality constraints. Therefore, such a task offloading problem is in the form of (1).
@@ III-A Problem Transformation
    There exists at least one optimum $\mathbf{y}^{\star}$ to problem (6).
@@ III-B Algorithm Design
Our proposed algorithm maintains four variables $\mathbf{y}^{k}\in\mathbb{R}^{N}$, $\mathbf{q}^{k}\in\mathbb{R}^{np}$, $\mathbf{u}^{k}\in\mathbb{R}^{n(m+p)}$, and $\mathbf{z}^{k}\in\mathbb{R}^{n(m+p)}$, where $\mathbf{y}^{k}$ is the primal variable, $\mathbf{q}^{k}$ and $\mathbf{u}^{k}$ are the dual variables associated with the inequality constraint and the equality constraint in (6), respectively, and $\mathbf{z}^{k}$ is the auxiliary variable that enables the update of $\mathbf{u}^{k}$.
@@
(a) $[P^{W}]_{ij}=[P^{H}]_{ij}=0$, $\forall i\in\mathcal{V}$, $\forall j\notin\mathcal{N}_{i}$; (b) $P^{W}$ and $P^{H}$ are symmetric and positive semidefinite; (c) $P^{W}\mathbf{1}_{n}=\mathbf{1}_{n}$, $\operatorname{Null}(P^{H})=\operatorname{span}(\mathbf{1}_{n})$; (d) $P^{W}+P^{H}\prec I_{n}$.
There are many options for $P^{W},P^{H}$ under Assumption 2. For example, we may let $P^{W}\!=\!\frac{I+P^{\prime}}{2}$ and $P^{H}\!=\!\frac{I-P^{\prime}}{2}$ for some $P^{\prime}\in\mathbb{R}^{n\times n}$ such that $[P^{\prime}]_{ij}=[P^{\prime}]_{ji}>0$$\forall\{i,j\}\in\mathcal{E}$, $\sum_{j\in\mathcal{N}_{i}}[P^{\prime}]_{ij}=1$, $[P^{\prime}]_{ii}>0$, and $[P^{\prime}]_{ij}=0$ otherwise [25].
@@
$\mathbf{y}^{k+1}$ $=\mathcal{P}_{Y}[\mathbf{y}^{k}-\gamma\cdot(\mathbf{d}^{k})^{T}],$ (7)
@@
In the primal update (7), $\mathbf{d}^{k}$ is the Jacobian matrix (or equivalently, $(\mathbf{d}^{k})^{T}$ is the gradient) of the following augmented-Lagrangian-like function $\mathcal{R}^{k}(\mathbf{y})$ at $\mathbf{y}^{k}$:
@@
$\mathbf{d}^{k}=$ $\frac{\partial\mathbf{f}(\mathbf{y}^{k})}{\partial\mathbf{y}}+(W\mathbf{u}^{k}-\frac{1}{\rho}\mathbf{z}^{k})^{T}\mathbf{B}+\frac{1}{\rho}(\mathbf{y}^{k})^{T}\mathbf{B}^{T}\mathbf{B}$
$+(\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k}))^{T}\frac{\partial\mathbf{G}(\mathbf{y}^{k})}{\partial\mathbf{y}}+\frac{1}{\rho}\mathbf{c}^{T}\mathbf{B},$ (12)
where $\frac{\partial\mathbf{f}(\mathbf{y}^{k})}{\partial\mathbf{y}}$ and $\frac{\partial\mathbf{G}(\mathbf{y}^{k})}{\partial\mathbf{y}}$ are the Jacobian matrices of $\mathbf{f}(\mathbf{y})$ and $\mathbf{G}(\mathbf{y})$ at $\mathbf{y}^{k}$, respectively. Clearly, the primal update (7) is a gradient projection operation aiming at the descent of the augmented-Lagrangian-like function $\mathcal{R}^{k}(\mathbf{y})$ in (11).
The dual update (8) adopts the idea of virtual queue in [36] (yet our primal update is completely different from that in [36]), which allows $\mathbf{q}^{k}$ to emulate the Lagrange multiplier associated with the inequality constraint $\mathbf{G}(\mathbf{y})\leq\mathbf{0}$ in (6) and benefits the convergence rate analysis compared to conventional dual subgradient updates.
@@
where $\mathbf{u}=[u_{1}^{T},\ldots,u_{n}^{T}]^{T}$. We first make an attempt to solve the dual problem (13) by employing the distributed first-order method P-EXTRA [42]. According to [37], such a dual P-EXTRA algorithm gives
@@
where $\mathbf{z}^{k}$ takes the same recursive form as (10) with $\mathbf{z}^{0}=\rho H\mathbf{u}^{0}$. Note that the above updates are not implementable because in general the dual function $D_{e}(\mathbf{u})$ is inaccessible.
@@
then $\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}$ is also a subgradient of $D_{e}(\mathbf{u}^{k+1})$ like $\mathbf{z}^{k}+\rho(\mathbf{u}^{k+1}-W\mathbf{u}^{k})$. Hence, we choose to equalize these two items, leading to the update equation of $\mathbf{u}^{k+1}$ which is exactly (9). Then, to derive a legitimate update equation for $\mathbf{y}^{k}$, note that (15) can be rewritten as $-\mathbf{B}^{T}\mathbf{u}^{k+1}\in\partial\mathbf{f}(\mathbf{y}^{k+1})$, which, along with (9), results in $-\mathbf{B}^{T}(W\mathbf{u}^{k}+\frac{1}{\rho}(\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}-\mathbf{z}^{k}))\in\partial\mathbf{f}(\mathbf{y}^{k+1})$. This is equivalent to
@@
where $\mathcal{R}^{k}(\mathbf{y})$ is given in (11) (with $\mathbf{G}(\mathbf{y})\equiv\mathbf{0}$). Therefore, (16), (9), and (10) together constitute an equivalent form of the aforementioned dual P-EXTRA algorithm, i.e., (14) and (10). Recall that (7) is a gradient projection operation for emulating (16). Consequently, for problem (6) with equality constraints only, our proposed algorithm can be viewed as a dual first-order method that originates from dual P-EXTRA yet introduces the gradient projection operation to acquire readily available and computationally efficient estimates of dual subgradients.
@@
Our proposed algorithm does not directly minimize $\mathcal{R}^{k}(\mathbf{y})$, i.e., adopt (16) as the primal update. This is because $f_{i}$ and $g_{i}$ appearing in $\mathcal{R}^{k}(\mathbf{y})$ are functions of coupled variables, which disable the decentralized realization of (16). Similarly, the primal update of IPLUX in [25] minimizes $\mathcal{R}^{k}(\mathbf{y})$ with $\mathbf{f}(\mathbf{y})$ substituted by its first-order approximation, which can only tackle problems without variable coupling and, thus, is unable to solve problem (1) without any centralized coordination. In contrast, the gradient of $\mathcal{R}^{k}(\mathbf{y})$ has a decomposable structure, so that the gradient projection step (7) can be executed decentralizedly, which will be elaborated in Section III-C. Also, (7) can significantly reduce the computational complexity of (16). Nevertheless, by adopting (7) instead of (16), we indeed compromise on the accuracy in estimating dual subgradients, thereby intensifying the challenge of convergence analysis.

### v4

@@ II-A Examples of Applications
Many real-world engineering problems can be cast into the form of (1). Below, we provide two such examples.
@@ II-A1 A Multi-Vehicle Formation Stabilization Problem
Consider a multi-vehicle formation stabilization problem. Our goal is to stabilize $n$ vehicles on a plane toward an equilibrium point in a decentralized way. For each vehicle $i$, $s_{i}^{k}=[(p_{i}^{k})^{T},(v_{i}^{k})^{T}]^{T}\in\mathbb{R}^{4}$ is the state and $u_{i}^{k}\in\mathbb{R}^{2}$ is the control input at time $k\geq 0$, where $p_{i}^{k}\in\mathbb{R}^{2}$, $v_{i}^{k}\in\mathbb{R}^{2}$ are the position and the velocity of vehicle $i$, respectively. The system dynamics of each vehicle $i$ is given by
@@
Given the initial states $s_{i}^{0}$ $\forall i=1,\ldots,n$, all the vehicles obtain their control actions through model predictive control (MPC), which requires them to cooperatively solve a constrained convex optimization problem. The global cost is $L(s,u)=\sum_{i=1}^{n}\sum_{k=0}^{N-1}\alpha_{1}\sum_{j\in\mathcal{N}_{i}}\|p_{i}^{k}-p_{j}^{k}+d_{ij}\|^{2}+\alpha_{2}\|v_{i}^{k}\|^{2}+\alpha_{3}\|u_{i}^{k}\|^{2}+L^{d}_{i}$, where $N\geq 1$ is the prediction horizon in MPC, $\alpha_{i}\forall i={1,2,3}$ are positive weighting constants, and $d_{ij}$ is the desired equilibrium position between node $i$ and $j$. The tracking cost $L^{d}_{i}$ is defined as $\alpha_{4}\|\sum_{i=1}^{m}p_{i}^{k}/m\!-\!p_{d}\|^{2}$ ($m<n$) if $i$ is one of the $m$ core vehicles; otherwise, $L^{d}_{i}=0$ [dunbar2006distributed]. Let the optimization variable associated with each vehicle $i$ be $x_{i}=[(s_{i}^{1})^{T},\ldots,(s_{i}^{N-1})^{T},(u_{i}^{0})^{T},\ldots,(u_{i}^{N-1})^{T}]^{T}\in\mathbb{R}^{6N}$. The multi-vehicle formation stabilization problem is formulated as the following optimization problem
@@
In (3), each $H_{i}$ is a symmetric positive semidefinite matrix and each $Q_{i}\in\mathbb{R}^{6|\mathcal{N}_{i}|N}$, which can be obtained from the definition of $x_{i}$ and $L(s,u)$. In (3), each vehicle $i$’s local cost is influenced by the positions of its neighboring vehicles [dunbar2006distributed]. It can be verified that problem (3) satisfies Assumption 1. The first constraint in (3) stands for resource limitations and position/velocity requirements across the whole system. The second constraint represents the dynamics of each vehicle $i$ over the prediction horizon, which is directly derived from (2). The last constraint restricts each $x_{i}$ to some compact convex set $X_{i}$ due to certain system limitations. In (3), $H_{i}$, $Q_{i}$, $P_{i}$, $b_{i}$, $C_{i}$, $b_{i}^{\prime}$, and $X_{i}$ are accessible only to vehicle $i$. Clearly, the distributed MPC problem (3) is a particular form of (1).
@@ II-A2 Task Offloading over Cooperative Mobile Edge Computing Networks
Consider a task offloading problem among $n$ edge servers forming a connected network. Each server receives a large number of latency-sensitive tasks from exogenous sources (e.g., mobile devices) and, due to limited and heterogeneous computation resources, needs to offload part of its tasks to neighboring servers to make full use of computation resources[xiao2018distributed]. All the edge servers cooperate to find optimal task offloading decisions by minimizing the total completion time of all the tasks subject to the limitations of channel bandwidth or computation capability.
Let $\mathcal{V}=\{1,2,\ldots,n\}$ denote the set of edge servers, and $\mathcal{N}_{i}$ the neighbor set of edge server $i$ including itself. Define $x_{ii}$ and $x_{ij}$ as the quantity of tasks allocated by edge server $i$ to itself and to its neighboring edge server $j$, respectively, and $x_{i}=[x_{ij}]_{\forall j\in\mathcal{N}_{i}}$ as the task offloading decision of edge server $i$. Consider an $M/M/1$ queuing system for each edge server to process its received tasks, where the task arrival $\lambda_{i}$ per unit time follows a Poisson distribution and $\mu_{i}$ is the maximum task processing rate [wang2024distributed]. According to the queuing theory, the processing latency of edge server $i$ is $f_{i}\left(x_{\mathcal{N}_{i}}\right)=\frac{\sum_{j\in\mathcal{N}_{i}}\lambda_{j}x_{ji}}{\mu_{i}-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}x_{ji}}$ and the computing capacity constraint is $g_{i}\left(x_{\mathcal{N}_{i}}\right)=\sum_{j\in\mathcal{N}_{i}}\lambda_{j}x_{ji}-\mu_{i}$, where both $f_{i}$ and $g_{i}$ are determined by the tasks offloaded by both edge server $i$ itself and its neighbors, i.e., $x_{ji}$ $\forall j\in\mathcal{N}_{i}$. For each edge server, the sum of the task offloaded to itself and neighboring servers should be equal to the tasks it received from the exogenous sources, leading to linear equality constraints. Therefore, such a task offloading problem in [wang2024distributed] is in the form of (1).
@@ III-A Problem Transformation
    There exists at least one optimum $\mathbf{y}^{\star}$ to problem (6). Any optimal solution $\mathbf{y}^{\star}=[(x_{1}^{\star})^{T},(t_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T},(t_{n}^{\star})^{T}]^{T}$ to problem (6) satisfies that $(x_{1}^{\star})^{T},\ldots,(x_{n}^{\star})^{T}]^{T}$ is an optimal solution to problem (1).
@@
Proposition 1(d) shows that problem (6) and problem (1) are equivalent, which allows us to obtain the optimal solution of problem (1) by solving problem (6).
@@ III-B Algorithm Design
Our proposed algorithm maintains four variables $\mathbf{y}^{k}\in\mathbb{R}^{N}$, $\mathbf{q}^{k}\in\mathbb{R}^{np}$, $\mathbf{u}^{k}\in\mathbb{R}^{n(m+p)}$, and $\mathbf{z}^{k}\in\mathbb{R}^{n(m+p)}$, where $\mathbf{y}^{k}$ is the primal variable, $\mathbf{q}^{k}$ and $\mathbf{u}^{k}$ are the dual variables associated with the inequality constraint and the equality constraint in (6), respectively, and $\mathbf{z}^{k}$ is an auxiliary variable whose role will be explained shortly.
@@
(a) $P^{W}]_{ij}=[P^{H}]_{ij}=0$, $\forall i\in\mathcal{V}$, $\forall j\notin\mathcal{N}_{i}$; (b) $P^{W}$ and $P^{H}$ are symmetric and positive semidefinite; (c) $P^{W}\mathbf{1}_{n}=\mathbf{1}_{n}$, $\operatorname{Null}(P^{H})=\operatorname{span}(\mathbf{1}_{n})$; (d) $P^{W}+P^{H}\preceq I_{n}$.
Assumption 2(a) is used to enable distributed implementation, while (b)–(d) are imposed to ensure convergence; see [wu2022distributed, ShiW2015a, wu2019improved] for detailed discussions. There are many options for $P^{W},P^{H}$ under Assumption 2. For example, we may let $P^{W}\!=\!\frac{I+P^{\prime}}{2}$ and $P^{H}\!=\!\frac{I-P^{\prime}}{2}$ for some $P^{\prime}\in\mathbb{R}^{n\times n}$ such that $P^{\prime}]_{ij}=[P^{\prime}]_{ji}>0$ $\forall\{i,j\}\in\mathcal{E}$, $\sum_{j\in\mathcal{N}_{i}}[P^{\prime}]_{ij}=1$, $P^{\prime}]_{ii}>0$, and $P^{\prime}]_{ij}=0$ otherwise [wu2022distributed].
@@
$\mathbf{y}^{k+1}$ $=\mathcal{P}_{Y}[\mathbf{y}^{k}-\gamma\cdot\mathbf{d}^{k}],$ (7)
@@
In the primal update (7), $\mathbf{d}^{k}$ is the gradient of the following augmented-Lagrangian-like function $\mathcal{R}^{k}(\mathbf{y})$ at $\mathbf{y}^{k}$:
@@
$\mathbf{d}^{k}=$ $\nabla\mathbf{f}(\mathbf{y}^{k})+\mathbf{B}^{T}(W\mathbf{u}^{k}-\frac{1}{\rho}\mathbf{z}^{k})+\frac{1}{\rho}\mathbf{B}^{T}\mathbf{B}\mathbf{y}^{k}$
$+(\frac{\partial\mathbf{G}(\mathbf{y}^{k})}{\partial\mathbf{y}})^{T}(\mathbf{q}^{k}+\mathbf{G}(\mathbf{y}^{k}))-\frac{1}{\rho}\mathbf{B}^{T}\mathbf{c},$ (12)
where $\frac{\partial\mathbf{G}(\mathbf{y}^{k})}{\partial\mathbf{y}}$ is the Jacobian matrix of $\mathbf{G}(\mathbf{y})$ at $\mathbf{y}^{k}$. Clearly, the primal update (7) is a gradient projection operation aiming at the descent of the augmented-Lagrangian-like function $\mathcal{R}^{k}(\mathbf{y})$ in (11).
The dual update (8) adopts the idea of virtual queue in [YuHao2017], which was originally developed for addressing inequality-constrained convex problems. It can be viewed as a modification of the dual subgradient algorithm with an additional proximal term in the primal update and an approximate dual update that resembles a queuing equation. In the absence of the equality constraints in (6), minimizing $\mathcal{R}^{k}(\mathbf{y})$ plus a proximal term $\frac{\gamma}{2}\|\mathbf{y}-\mathbf{y}^{k}\|^{2}$ corresponds to the primal update of the virtual queue algorithm in [YuHao2017], while (8) matches its dual update. However, this primal update cannot be implemented in a distributed manner because of the variable coupling in the local functions. This motivates us to employ a different primal update (7) for distributed implementation, and retain the dual update (8) for effectively handling the inequality constraints.
@@
where $\mathbf{u}=[u_{1}^{T},\ldots,u_{n}^{T}]^{T}$. We first make an attempt to solve the dual problem (13) by employing the distributed first-order method P-EXTRA [ShiW2015a]. According to [wu2019improved], such a dual P-EXTRA algorithm gives
@@
where $\mathbf{z}^{k}$ takes the same recursive form as (10) with $\mathbf{z}^{0}=\rho H\mathbf{u}^{0}$. Here, $\mathbf{z}^{k}$ is the dual variable associated with the consensus constraint $u_{1}=\ldots=u_{n}$ in problem (13). According to the update equation (10) with $\mathbf{z}^{0}=\rho H\mathbf{u}^{0}$, we have $\mathbf{z}^{k}=\rho\sum_{\ell=0}^{k}H\mathbf{u}^{\ell}$, which is the accumulation of the consensus error over iterations and facilitates the consensus of all the $u_{i}^{\ell}$’s. Note that the above updates are not implementable because in general the dual function $D_{e}(\mathbf{u})$ is inaccessible.
@@
then $\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}$ is also a subgradient of $D_{e}(\mathbf{u}^{k+1})$ like $\mathbf{z}^{k}+\rho(\mathbf{u}^{k+1}-W\mathbf{u}^{k})$. Hence, we choose to equalize these two items, leading to the update equation of $\mathbf{u}^{k+1}$ which is exactly (9). Then, to derive a legitimate update equation for $\mathbf{y}^{k}$, note that (15) can be rewritten as $\mathbf{B}^{T}\mathbf{u}^{k+1}\in\nabla\mathbf{f}(\mathbf{y}^{k+1})$, which, along with (9), results in $\mathbf{B}^{T}(W\mathbf{u}^{k}+\frac{1}{\rho}(\mathbf{B}\mathbf{y}^{k+1}-\mathbf{c}-\mathbf{z}^{k}))\in\nabla\mathbf{f}(\mathbf{y}^{k+1})$. This is equivalent to
@@
where $\mathcal{R}^{k}(\mathbf{y})$ is given in (11) (with $\mathbf{G}(\mathbf{y})\equiv\mathbf{0}$). Therefore, (16), (9), and (10) together constitute an equivalent form of the aforementioned dual P-EXTRA algorithm, i.e., (14) and (10). Recall that (7) is a gradient projection operation for emulating (16). Consequently, for problem (6) without the inequality constraint, our proposed algorithm can be viewed as a dual first-order method that originates from dual P-EXTRA yet introduces the gradient projection operation to acquire readily available and computationally efficient estimates of dual subgradients.
@@
Our proposed algorithm does not directly minimize $\mathcal{R}^{k}(\mathbf{y})$, i.e., adopt (16) as the primal update. This is because $f_{i}$ and $g_{i}$ appearing in $\mathcal{R}^{k}(\mathbf{y})$ are functions of coupled variables, which disable the decentralized realization of (16). Similarly, the primal update of IPLUX in [wu2022distributed] minimizes $\mathcal{R}^{k}(\mathbf{y})$ with $\mathbf{f}(\mathbf{y})$ substituted by its first-order approximation, which can only tackle problems without variable coupling and, thus, is unable to solve problem (1) without any centralized coordination. In contrast, the gradient of $\mathcal{R}^{k}(\mathbf{y})$ has a decomposable structure, so that the gradient projection step (7) can be executed decentralizedly, which will be elaborated in Section III-C. Also, (7) can significantly reduce the computational complexity of (16). Nevertheless, by adopting (7) instead of (16), we indeed compromise on the accuracy in estimating dual subgradients, thereby intensifying the challenge of convergence analysis.